(20%) A production operation is making 150 units of a product by engaging five
workers for 300 hours. However, 40 percent of the units appear to have various quality
problems, and the company decides to sell them as seconds at a price of £50 each
when a normal unit is sold for £150. To improve the situation, several initiatives are
proposed, including a scheme where, for every improvement, 50 percent will be given
to workers and the other 50 percent will be held by the company. This results in a
significant drop in defects as now only 10 units are faulty out of an output of 130 units. Determine the appropriate bonus per hour for the workers under the bonus
scheme if the cost per piece is £70 both before and after the scheme. (10%)
Each worker is paid for 300 hours of labour
Amount workers receive from improvements
1
Expert's answer
2019-12-16T10:18:49-0500

The amount of normal units increases from 150×(1 - 0.4) = 90 to 130 - 10 = 120, so all the workers will receive 30×(150 - 70)/2 = 1200, which is 1200/300/5 = £0.8 per unit.